# Semantically-correlated memories in a dense associative model

### v2

@@ 1.2 Motivations from neuroscience
A classical result in the hetero-association neuroscience literature is due to Miyashita (1988). This work demonstrated hetero-association of stimuli in monkey temporal cortex could arise semantically via repeated presentations of the same stimuli in the same order, not only spatially via similarities in the stimuli themselves. Miyashita (1988) showed neurons responsive to presentation of randomly-generated fractal patterns had a monotonically-decreasing auto-correlation between the firing rates due to the current pattern and the next expected patterns, up to a distance of 6 patterns into the future.
Work on numerosity in birds, non-human primates, and humans (Nieder et al., 2002; Ditz & Nieder, 2015; Nieder, 2012; Kutter et al., 2018) have repeatedly provided evidence of neurons responding to specific numbers or quantities. In these experiments, the stimuli (numbers or quantities) can be both semantically and spatially correlated – i.e., they can have the known semantic ordering of $1,2,3\dots$ or ‘some, more, even more …’, as well as the spatial or statistical relationships between the stimuli. Notably, even in abstract number experiments where spatial correlations are moot, semantic distances up to a range of $\sim 5$ numbers [^3] Depending on the species, brain area, and stimuli modality. (as measured by significant auto-correlations of the neural activity) are common.
@@ 1.3 Motivations from machine learning
In Appendix A.1, I summarise the technical connection between Transformers and associative memory, and how Transformers’ attention mechanisms take the hetero-associative form mathematically. Functionally, however, the attention mechanism is not obliged to perform hetero-association, since its values and keys are created independently by their respective weight matrices (see Vaswani et al. (2017)) and can in-principle align these functionally so as to perform auto-association, or otherwise some mixture of auto- and hetero-association. Taking this perspective seriously opens the way for analysing Transformers through the lens of potential mixtures of auto- and hetero-associative dynamics, à la the analysis of a large language model in Ramsauer et al. (2021) by considering the implied energy landscapes in each of its attention heads. For this to be possible, however, a first step is to rigorously develop and study a dense auto- and hetero-association model and its inherent computational capabilities. (In Appendix A.2, I provide suggestions for new interpretation approaches based on this paper’s results.)
1.4 Our contributions
@@ 2.2 Theoretical analysis
A typical analysis to perform on associative memory networks is to probe its memory storage capacity, i.e., how many memories can be stored given $n$ neurons? In CDAM, when $a,h\neq 0$, the regular notions of ‘capacity’ seem inapplicable. This is because ‘capacity’ is normally measured in the pure auto-associative case by giving a noise-corrupted or partial memory pattern, and observing whether and how closely the model’s dynamics converge to the uncorrupted or complete memory pattern (e.g., see Amit et al. (1985) for the classical model and Demircigil et al. (2017) for the dense model). In the pure hetero-associative case, ‘capacity’ has (to my knowledge) only ever been studied in the linear sequences case (e.g., see Löwe (1998) for the classical model and Chaudhry et al. (2023) for the dense model). However, in this model I study general mixtures of both auto- and hetero-association, as well as arbitrary memory graphs (not just linear cycles). It is therefore unclear what an appropriate notion of ‘capacity’ for this mixture would be and what it would measure [^4] From a practical standpoint, however, a variant of auto-association wherein we care less about the absolute overlap between the pattern and the state and instead about which pattern has the highest overlap is possible, and something I explore in Appendix A.3..
One can, nevertheless, study the model in a similar spirit of analysis. To this end, we demonstrate the dynamics of our model in the thermodynamic limit. First, let us set aside the choice of $\eta$ which controls the amplitude of each step’s update. We define the overlap between a memory pattern and state as $m^{\mu}(t)=\frac{\sigma(t)^{T}\xi^{\mu}}{n}$. For an undirected memory graph $\mathcal{M}$ without loops, our energy function is
@@
In Appendix A.4, I show numerical simulations to demonstrate these four behavioural modes: auto-association, narrow hetero-association, wide hetero-association, and neutral quiescence. These simulations also demonstrate how we need $h=1-a$ (or $a=1-h$) for the mean neural activity to converge to 00 in the limit of $t\rightarrow\infty$ and $n\rightarrow\infty$ when we have random memory patterns, i.e., to keep an unbiased excitatory–inhibitory (E–I) balance [^5] Note this does not imply there is no activity in the network, since we allow neurons to take negative values..
@@
Of natural interest is when $h\neq 0$, which provides interactions between the patterns. What is interesting about Equation 1 is the possibility of both auto- and hetero-associative terms affecting the dynamics when both $a,h\neq 0$. As implied informally above, this means the model cannot perform pure pattern retrieval, i.e., retrieval of a single memory pattern $\xi^{\mu}$ without at least partial retrieval of other patterns. To show this, it is useful to refer to the overlap between a pattern $\xi^{\mu}$ and a state $\sigma(t)$. For this, we use the Pearson product-moment correlation coefficient, which for pattern $\mu$ at time $t$ we denote $r(\mu^{(t)})$.
@@
Hebbian auto- and hetero-associative mixtures cannot perform pure pattern retrieval for patterns not isolated in the memory graph. Suppose $\mu$ is not an isolated vertex in $\mathcal{M}$. Let $a,h>0$. Then the model cannot perform pure pattern retrieval of $\xi^{\mu}$. To see this, let $\{\xi^{\mu},\xi^{v}\}\in\mathcal{E}$ if $\mathcal{M}$ is undirected, and let $(\xi^{\mu},\xi^{v})\in\mathcal{E}$ if $\mathcal{M}$ is directed. Setting $\sigma(t)=\xi^{\mu}$ will cause the second term of $Q$ to be non-negative because $h>0$, and therefore $r(v^{(t+1)})$ will be proportionally large. Simultaneously, $r(\mu^{(t+1)})$ will be non-vanishing, since $a>0$. Therefore, no non-isolated pattern can be purely retrieved.
@@
then the model can perform pure pattern retrieval of some memory patterns. The excitatory auto-associative result, where $a>0$ and $h=0$, is simply a weighted version of Theorems 1–3 from Ramsauer et al. (2021). The excitatory hetero-associative result, where $a=0$ and $h>0$, is indicated by the limited pure retrieval case above, with the added restriction that there exists only one memory pattern, $\xi^{v}$, projecting from $\xi^{\mu}$ in $\mathcal{M}$. This restriction is necessary because if the out-degree of $\xi^{\mu}$ was 00, then after setting $\sigma(t)=\xi^{\mu}$, the projection matrix $Q$ would be filled with zeros since $a=0$. Therefore, values of $\sigma$ would converge to a value of $-\tilde{\xi}$. If the out-degree of $\xi^{\mu}$ was $>1$, we would have the situation limited pure retrieval case, only with multiple memory patterns with large $r$ values (with their strengths proportional to the weights of their respective in-edges from $\xi^{\mu}$ in $\mathcal{M}$). Finally, we need to achieve $\sigma(t+1)=\xi^{v}$ (or something arbitrarily close) to have pure pattern retrieval of $\xi^{v}$, since $a=0$ means we will not have the luxury of additional time-steps to achieve convergence. Fortunately, by Theorem 4 of Ramsauer et al. (2021), we can get arbitrarily close by requiring sufficiently large values of $\beta$ and $\eta$ to update the state to $\xi^{v}$ in a single step.
@@ Acknowledgements
Thank you to anonymous reviewers for their constructive feedback. Thank you to Dima Krotov, Dia Taha, Mengsen Zhang, Horacio Rotstein, Vasiliki Liontou, Robert Tang, Dan Murfet, Adam Shai, Christopher Earls, Tomoki Fukai, Tatsuya Haga, participants in the ‘Math + Neuroscience’ program at the Institute for Computational and Experimental Research in Mathematics, participants in the ‘Physics of Life’ symposium in November 2023 at the Center for the Physics of Biological Function, and many others for thoughtful comments and discussions.
@@ A.1 Connection between associative memory and Transformers
with $\xi_{i}^{\mu}$ (=±$=\pm 1$) static variables being the $p$ binary memory pattern vectors stored, and where $\xi_{\delta}^{\mu}$ is the product of the $\delta$-indexed variables in $\xi^{\mu}$.
@@
$\sum_{j=1}^{n}w_{ij}s_{j}^{(t)},\hskip 4.33382pt\text{where}\hskip 4.33382ptw_{ij}=\frac{1}{n}\sum_{\mu=1}^{p}\xi_{i}^{\mu}\xi_{j}^{\mu}.$
@@
Those familiar with Transformers (Vaswani et al., 2017) from the machine learning literature will recognise that Equation 5 is very close that of the attention mechanism. The only difference between Equation 5 and the attention mechanism is that one of the $\Xi$ variables should be replaced by a unique matrix. This connection was noticed by Ramsauer et al. (2021), and as Section 3.5 of Millidge et al. (2022) notes, associative memory models of this kind perform auto-association ($\sigma$ dynamically moves towards the memory patterns stored in $\Xi$) whereas the attention mechanism of Transformers performs hetero-association ($\sigma$ dynamically moves towards columns of a matrix $\Xi^{\prime}$ based on the corresponding closeness to columns in a separate matrix $\Xi$). However, as I note in the current work, the Transformer creates $\Xi$ and $\Xi^{\prime}$ from distinct weight matrices and there is nothing preventing those weight matrices from being partially or even entirely equivalent. The consequence of this is that Transformers, while their mathematical form is of a hetero-associative nature, can in fact implement any mixture of auto- and hetero-association.
@@ A.3 Non-traditional auto-association performance
First, it bears repeating that if we set $a=1$ and $h=0$, CDAM becomes the model of Equation 13 in Lucibello & Mézard (2023), where scaling comes from $a$. Analysing the performance of auto-associative recall is then done in the usual way of setting $\sigma$ as a noisy version of a memory pattern, then running the dynamics forward until convergence, and measuring the final overlap between the original (denoised) memory pattern and the final, convergent state. From a practical standpoint, however, it is possible to study the following non-traditional variation: we no longer not care about the absolute overlap between the pattern and the state, instead we simply care about which pattern has the highest overlap.
@@
As in MESH, I test CDAM on the FashionMNIST dataset (Xiao et al., 2017), which consists of $28\times 28$ grayscale images associated with a label from 10 classes of clothing (shirts, pants, shoes, etc.). We arrange the maximum-normalised pixel values of these images into vectors and use them as our memory states, i.e., $n=784=28\times 28$. As in previous simulations, I use $\beta=1$, and $\eta=0.1$. To initialise the network state, we choose a memory pattern $\mu$ (an image from the FashionMNIST dataset) and set $\sigma(0)=\xi^{\mu}+c\zeta$, where $\zeta$ is a random vector with elements independently drawn from the interval $[-0.5,0.5]$ and $c\in\mathbb{R}^{+}$ is the amplitude of the additive random noise. Here I use $c=1$. I run the simulation until convergence and measure the overlap of the final state with all memory vectors. Whichever memory vector has the highest overlap is considered to be ‘predicted’ (see Figure 8).
Figure 8: Examples of two FashionMNIST memory items in original (left), noisy (centre), and ‘predicted’ (right) forms.
@@ A.5 Biological mechanisms by which to learn $M$
In this context, it does not seem far-fetched to imagine a biological organism explicitly learning arbitrary hetero-associative structures of the form $M$, in CDAM, or some approximation thereof. This would most easily be achieved through repeated exposure to sequences of memory items as a random walk on $\mathcal{M}$, as was studied in human participants by Schapiro et al. (2013), who indeed found the participants learnt the community structure of the $\mathcal{M}$ equivalent in their study. However, it is quite likely animals encounter stimuli which have similar structure, and thereby use past experience as a cognitive bias to more efficiently store related structures and learn new instances of similar structures. A computational example of this idea comes from Whittington et al. (2020), who introduced the ‘Tolman-Eichenbaum Machine’ (TEM), a joint where-what model of hippocampus and entorhinal cortex which demonstrates close correspondence to biological data as well as having a mathematical relationship to Transformers (Whittington et al., 2022). This has been followed by alluring biological studies (El-Gaby et al., 2023), further solidifying the potential biological bases of learning and leveraging such structures and biases.
@@ A.7 Replication of Miyashita (1988)
As described in Subsection 1.2, Miyashita (1988) is a a classical study in the semantic hetero-association neuroscience literature, which showed neurons from monkey temporal cortex were responsive to the presentation of stimuli according to the order in which they were presented. These semantic links were developed without general regard to any spatial or statistical similarities shared between the stimuli. In neurons which were significantly responsive to the stimuli, their activity was significantly auto-correlated with the activity elicited by the stimuli up to a distance of 6 patterns into the future.

### v3

@@ 1.2 Motivations from neuroscience
A classical result in the hetero-association neuroscience literature is due to Miyashita (1988). This work demonstrated hetero-association of stimuli in monkey temporal cortex could arise semantically via repeated presentations of the same stimuli in the same order, not only spatially via similarities in the stimuli themselves. Miyashita (1988) showed neurons responsive to presentation of randomly-generated fractal patterns had a monotonically-decreasing auto-correlation between the firing rates due to the current pattern and the next expected patterns, up to a distance of 6 patterns into the past or future of the stimuli sequence.
Work on numerosity in birds, non-human primates, and humans (Nieder et al., 2002; Ditz & Nieder, 2015; Nieder, 2012; Kutter et al., 2018) have repeatedly provided evidence of neurons responding to specific numbers or quantities. In these experiments, the stimuli (numbers or quantities) can be both semantically and spatially correlated – i.e., they can have the known semantic ordering of ‘$1,2,3\dots$’ or ‘some, more, even more …’, as well as the spatial or statistical relationships between the stimuli, e.g., visually, the numerals ‘4’ and ‘9’. Notably, even in abstract number experiments where spatial correlations are moot, semantic distances up to a range of $\sim 5$ numbers [^3] Depending on the species, brain area, and stimuli modality. (as measured by significant auto-correlations of the neural activity) are common.
@@ 1.3 Motivations from machine learning
In Appendix A.1, I summarise the technical connection between Transformers and associative memory, and how Transformers’ attention mechanisms take the hetero-associative form mathematically. Functionally, however, the attention mechanism is not obliged to perform hetero-association, since its values and keys are created by their distinct weight matrices (see Vaswani et al. (2017)) and can in-principle align these functionally so as to perform auto-association, or otherwise some mixture of auto- and hetero-association. Taking this perspective seriously opens the way for analysing Transformers through the lens of potential mixtures of auto- and hetero-associative dynamics, à la the analysis of a large language model in Ramsauer et al. (2021) by considering the implied energy landscapes in each of its attention heads. For this to be possible, however, a first step is to rigorously develop and study a dense auto- and hetero-association model and its inherent computational capabilities. (In Appendix A.2, I provide suggestions for new interpretation approaches based on this paper’s results.)
1.4 Contributions
@@ 2.2 Theoretical analysis
A typical analysis to perform on associative memory networks is to probe its memory storage capacity, i.e., how many memories can be stored and reliably retrieved given $n$ neurons? In CDAM, when $a,h\neq 0$, the regular notions of ‘capacity’ seem inapplicable. This is because ‘capacity’ is normally measured in the pure auto-associative case by giving a noise-corrupted or partial memory pattern, and observing whether and how closely the model’s dynamics converge to the uncorrupted or complete memory pattern (e.g., see Amit et al. (1985) for the classical model and Demircigil et al. (2017) for the dense model). In the pure hetero-associative case, ‘capacity’ has (to my knowledge) only ever been studied in the linear sequences case (e.g., see Löwe (1998) for the classical model and Chaudhry et al. (2023) for the dense model). However, in this model I study general mixtures of both auto- and hetero-association, as well as arbitrary memory graphs (not just linear cycles). It is therefore unclear what an appropriate notion of ‘capacity’ for this mixture would be and what it would measure [^4] From a practical standpoint, however, a non-traditional variant of auto-association wherein we care less about the absolute overlap between the pattern and the state and instead about which pattern has the highest overlap is possible, and something I explore in Appendix A.3..
One can, nevertheless, study the model in a similar spirit of analysis. To this end, I demonstrate the dynamics of the model in the thermodynamic limit. First, let us set aside the choice of $\eta$ which controls the amplitude of each step’s update. We define the overlap between a memory pattern and state as $m^{\mu}(t)=\frac{\sigma(t)^{T}\xi^{\mu}}{n}$. For an undirected memory graph $\mathcal{M}$ without loops, the energy function is
@@
In Appendix A.4, I show numerical simulations to demonstrate these four behavioural modes: auto-association, narrow hetero-association, wide hetero-association, and neutral quiescence. These simulations also demonstrate how we need $a+h=1$ for the mean neural activity to converge to 00 in the limit of $t\rightarrow\infty$ and $n\rightarrow\infty$ when we have random memory patterns, i.e., to keep an unbiased excitatory–inhibitory (E–I) balance [^5] Note this does not imply there is no activity in the network, since we allow neurons to take negative values..
@@
Of natural interest is when $h\neq 0$, which provides interactions between the patterns. What is interesting about Equation 1 is the possibility of both auto- and hetero-associative terms affecting the dynamics when both $a,h\neq 0$. As implied informally above, this means the model cannot perform pure pattern retrieval, i.e., retrieval of a single memory pattern $\xi^{\mu}$ without at least partial retrieval of other patterns. To show this, it is useful to refer to the overlap between a pattern $\xi^{\mu}$ and a state $\sigma(t)$. For this, we can use the Pearson product-moment correlation coefficient, which for pattern $\mu$ at time $t$ I denote $r(\mu^{(t)})$.
@@
Hebbian auto- and hetero-associative mixtures cannot perform pure pattern retrieval for patterns not isolated in the memory graph. Suppose $\mu$ is not an isolated vertex in $\mathcal{M}$. Let $a,h>0$. Then the model cannot perform pure pattern retrieval of $\xi^{\mu}$. To see this, let $\{\xi^{\mu},\xi^{v}\}\in\mathcal{E}$ if $\mathcal{M}$ is undirected, and let $(\xi^{\mu},\xi^{v})\in\mathcal{E}$ if ${\bm{v}}{\mathcal{M}}$ is directed. Setting $\sigma(t)=\xi^{\mu}$ will cause the second term of $Q$ to be non-negative because $h>0$, and therefore $r(v^{(t+1)})$ will be proportionally large. Simultaneously, $r(\mu^{(t+1)})$ will be non-vanishing, since $a>0$. Therefore, no non-isolated pattern can be purely retrieved.
@@
then the model can perform pure pattern retrieval of some memory patterns. The excitatory auto-associative result, where $a>0$ and $h=0$, is simply a weighted version of Theorems 1–3 from Ramsauer et al. (2021). The excitatory hetero-associative result, where $a=0$ and $h>0$, is indicated by the limited pure retrieval case above, with the added restriction that there exists only one memory pattern, $\xi^{v}$, projecting from $\xi^{\mu}$ in $\mathcal{M}$. This restriction is necessary because if the out-degree of $\xi^{\mu}$ was 00, then after setting $\sigma(t)=\xi^{\mu}$, the projection matrix $Q$ would be filled with zeros since $a=0$. Therefore, values of $\sigma$ would converge to a value of $-\tilde{\xi}$. If the out-degree of $\xi^{\mu}$ was $>1$, we would have a limited pure retrieval case, only with multiple memory patterns with large $r$ values (with their strengths proportional to the weights of their respective in-edges from $\xi^{\mu}$ in $\mathcal{M}$). Finally, we need to achieve $\sigma(t+1)=\xi^{v}$ (or something arbitrarily close) to have pure pattern retrieval of $\xi^{v}$, since $a=0$ means we will not have the luxury of additional time-steps to achieve convergence. Fortunately, by Theorem 4 of Ramsauer et al. (2021), we can get arbitrarily close by requiring sufficiently large values of $\beta$ and $\eta$ to update the state to $\xi^{v}$ in a single step.
@@ Acknowledgements
Thank you to anonymous reviewers for their constructive feedback. Thank you to Dima Krotov, Mengsen Zhang, Horacio Rotstein, Vasiliki Liontou, Robert Tang, Dan Murfet, Adam Shai, Dia Taha, Christopher Earls, Tomoki Fukai, Tatsuya Haga, participants in the ‘Math + Neuroscience’ program at the Institute for Computational and Experimental Research in Mathematics, participants in the ‘Physics of Life’ symposium in November 2023 at the Center for the Physics of Biological Function, and many others for thoughtful comments and discussions.
@@ A.1 Connection between associative memory and Transformers
with $\xi_{i}^{\mu}$ (=±$=\pm 1$) static variables being the $p$ binary memory pattern vectors stored, and where $\xi_{\delta}^{\mu}$ is the product of the $\delta$-indexed components in $\xi^{\mu}$.
@@
$\sum_{j=1}^{n}w_{ij}\sigma_{j}^{(t)},\hskip 4.33382pt\text{where}\hskip 4.33382ptw_{ij}=\frac{1}{n}\sum_{\mu=1}^{p}\xi_{i}^{\mu}\xi_{j}^{\mu}.$
@@
Those familiar with Transformers (Vaswani et al., 2017) from the machine learning literature will recognise that Equation 5 is very close that of the attention mechanism. A difference between Equation 5 and the attention mechanism is that one of the $\Xi$ variables should be replaced by a unique matrix. This connection was noticed by Ramsauer et al. (2021), and as Section 3.5 of Millidge et al. (2022) notes, associative memory models of this kind perform auto-association ($\sigma$ dynamically moves towards the memory patterns stored in $\Xi$) whereas the attention mechanism of Transformers performs hetero-association ($\sigma$ dynamically moves towards columns of a matrix $\Xi^{\prime}$ based on the corresponding closeness to columns in a separate matrix $\Xi$). However, as I note in the current work, the Transformer creates $\Xi$ and $\Xi^{\prime}$ from distinct weight matrices and there is nothing preventing those weight matrices from being partially or even entirely equivalent. The consequence of this is that Transformers, while their mathematical form is of a hetero-associative nature, can in fact implement any mixture of auto- and hetero-association.
@@ A.3 Non-traditional auto-association performance
First, it bears repeating that if we set $a=1$ and $h=0$, CDAM becomes the model of Equation 13 in Lucibello & Mézard (2023), where scaling comes from $a$. Analysing the performance of auto-associative recall is then done in the usual way of setting $\sigma$ as a noisy version of a memory pattern, then running the dynamics forward until convergence, and measuring the final overlap between the original (denoised) memory pattern and the final, convergent state. From a practical standpoint, however, it is possible to study the following non-traditional variation: we no longer care about the absolute overlap between the pattern and the state, instead we simply care about which pattern has the highest overlap.
@@
As in MESH, I test CDAM on the FashionMNIST dataset (Xiao et al., 2017), which consists of $28\times 28$ grayscale images associated with a label from 10 classes of clothing (shirts, pants, shoes, etc.). I arrange the maximum-normalised pixel values of these images into vectors and use them as the memory states, i.e., $n=784=28\times 28$. As in previous simulations, I use $\beta=1$, and $\eta=0.1$. To initialise the network state, I choose a memory pattern $\mu$ (an image from the FashionMNIST dataset) and set $\sigma(0)=\xi^{\mu}+c\zeta$, where $\zeta$ is a random vector with elements independently drawn from the interval $[-0.5,0.5]$ and $c\in\mathbb{R}^{+}$ is the amplitude of the additive random noise. Here I use $c=1$. I run the simulation until convergence and measure the overlap of the final state with all memory vectors. Whichever memory vector has the highest overlap is considered to be ‘predicted’ (see Figure 8).
Figure 8: Examples of two FashionMNIST memory items in original (left), noisy (center), and ‘predicted’ (right) forms.
@@ A.5 Biological mechanisms by which to learn $M$
In this context, it does not seem far-fetched to imagine a biological organism explicitly learning arbitrary hetero-associative structures of the form $M$, in CDAM, or some approximation thereof. This would most easily be achieved through repeated exposure to sequences of memory items as a random walk on $\mathcal{M}$, as was studied in human participants by Schapiro et al. (2013), who indeed found the participants learnt the community structure of the $\mathcal{M}$ equivalent in that study. However, it is quite likely animals encounter stimuli which have similar structure, and thereby use past experience as a cognitive bias to more efficiently store related structures and learn new instances of similar structures. A computational example of this idea comes from Whittington et al. (2020), who introduced the ‘Tolman-Eichenbaum Machine’ (TEM), a joint where-what model of hippocampus and entorhinal cortex which demonstrates close correspondence to biological data as well as having a mathematical relationship to Transformers (Whittington et al., 2022). This has been followed by alluring biological studies (El-Gaby et al., 2023), further solidifying the potential biological bases of learning and leveraging such structures and biases.
@@ A.7 Replication of Miyashita (1988)
As described in Subsection 1.2, Miyashita (1988) is a a classical study in the semantic hetero-association neuroscience literature, which showed neurons from monkey temporal cortex were responsive to the presentation of stimuli according to the order in which they were presented. These semantic links were developed without general regard to any spatial or statistical similarities shared between the stimuli. In neurons which were significantly responsive to the stimuli, their activity was significantly auto-correlated with the activity elicited by the stimuli up to a distance of 6 patterns into the past or future of the stimuli sequence.